OpenStudy (cory24):

What is the equation, in slope-intercept form, of the line that passes through (0, 3) and has a slope of −6?

OpenStudy (mathmale):

You are given a point (with x- and y-coordinates) and the slope of a straight line. Substitute this numeric data into the "point-slope formula for the equation of a straight line," and then re-write your result in the form y = mx + b.

OpenStudy (mathmale):

Alternatively, start with the "slope-intercept form" y = mx + b and substitute the given numeric info with the aim of finding the value of b.

OpenStudy (ajspeller):

(0,3) ==> m=-6 y=mx+b x=0 y=3 m=-6 make the substitutions and solve for b. then rewrite, y=-6x+b, where b is the value you just found above.